for the following equation, state the value of the discriminant and then describe the nature of the solutions. 4x^2-9x-1=0.
Value of the discriminant - (I got 97 - is that right?)
Equation has:
a) 2 imaginary solutions
b) 2 real solutions
c) 1 real solution
: for the following equation, state the value of the discriminant and then describe the nature of the solutions. 4x^2-9x-1=0.
Value of the discriminant - (I got 97 - is that right?)
Equation has:
a) 2 imaginary solutions
b) 2 real solutions
c) 1 real solution

Answer by Fombitz(1798)

b^2-4ac=(-9)^2-4(4)(-1)=81+16=97
Correct.
So what's the answer?? a,b, or c.